OpenStudy (anonymous):

Solve sqrt(x^5*y^19*z^10)

OpenStudy (anonymous):

please :)

OpenStudy (anonymous):

\[\text{Did you mean reduce?}\\ \sqrt{x^5y^{19}z^{10}}=[x^5y^{19}z^{10}]^\frac12=x^\frac52y^\frac{19}2z^5 \]

OpenStudy (anonymous):

Yes, my teacher has this as the solution: \[x^2y^9z^5\sqrt(xy)\] i'm not sure how he came up with it though